OpenStudy (anonymous):

oh God, I hate these questions: A coffee company want a new flavor of Cajun coffee. how many pounds of coffee worth $8 a pound should be added to 30 pounds of coffee worth $4 a pound to get a mixture worth $6 a pound?

OpenStudy (anonymous):

30 pounds.

OpenStudy (anonymous):

adds up to 60 pounds

OpenStudy (amistre64):

just to let you know, the term "oh god" is considered offensive by alot of people and should be avoided when used in this type of open forum please

OpenStudy (anonymous):

do you get it?

OpenStudy (anonymous):

amistre can you answer my question please? its over there!

OpenStudy (amistre64):

how many pounds of coffee worth $8 30 pounds of coffee worth $4 $6 a pound? amount@8 = 30(4-6)/(8-6) and cut off any negative signs

OpenStudy (amistre64):

ill lool in a minute

OpenStudy (anonymous):

thx

OpenStudy (amistre64):

30 looks good to me

OpenStudy (anonymous):

30lbs@$4=$120 you're adding X lbs@$8=$8X cost Total cost for mix = $120 + $8X Price/lb of mix = (120+8X)/(30+X) = $6 120+8X = 6(30+X) 2X = 60 X = 30lbs

OpenStudy (anonymous):

whoawhoawhoa! you did it the smart way!!!

OpenStudy (anonymous):

Well thank you!

OpenStudy (anonymous):

yup! it was all smartical :P